OpenStudy (anonymous):

Find the area of the region enclosed between the curve y=12x^2 + 30x and the x-axis.

OpenStudy (anonymous):

i got this, doing the same thing fam

OpenStudy (anonymous):

first you find the point of intersection

OpenStudy (anonymous):

My answer to the problem was -625/4 but the book says its 125/4.

OpenStudy (anonymous):

maybe your top curve was wrong

OpenStudy (anonymous):

Let me type down from where I think I made an error.

OpenStudy (anonymous):

\[\int\limits_{0}^{-5/2} 12x ^{2}+30x dx\]

OpenStudy (anonymous):

\[\left[ 4x ^{3}+15x ^{2} \right]_{0}^{-5/2} +k\]

OpenStudy (anonymous):

wrong way

OpenStudy (anonymous):

0 on the top and -5/2 on bottom

OpenStudy (anonymous):

OK, but I still end up with the same answer because when I evaluate the definite integral I'm still going to get a negative answer.

OpenStudy (anonymous):

i got positive

OpenStudy (anonymous):

125/4

OpenStudy (anonymous):

Can you show your work plox

OpenStudy (anonymous):

12x^2+30x= 0 so the intersection points are x = 0 and -5/2

OpenStudy (anonymous):

rite

OpenStudy (anonymous):

top curve is 12x^2+30x because you evaluate a number between 0 and -5/2 and it is greater than 0

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

same

OpenStudy (anonymous):

yea its negative number

OpenStudy (anonymous):

-156.25 .....

OpenStudy (anonymous):

I know to make it positive I need to swap a subtraction sign for and addition sign but I don't the proper procedure.

myininaya (myininaya):

\[12x^2+30x=0 \\ x=0 \text{ or } 12x+30=0 \\ x=0 \text{ or } x=\frac{-30}{12} \\ x=0 \text{ or } x=\frac{-5}{2} \] your limits are right upper limit=0 lower limit=-5/2 now we need to figure out which function is greater y=12x^2+30x or y=0 well you could plug in a number between the intersections to see.(Or you look at graph) I will do it the non-calculator way. A number between -5/2 and 0 is -1. y(x)=12x^2+30x y(-1)=12(-1)^2+30(-1)=12-30=-18 and y(x)=0 y(0)=0 so since -18<0 then 12x^2+30x<0 on the interval (-5/2,0) so y=0 is the greater function (The one above the other curve) \[\int\limits_\frac{-5}{2}^0 (0-[12x^2+30x]) dx\]

myininaya (myininaya):

\[\int\limits_\frac{-5}{2}^0 (-12x^2-30x) dx\]

myininaya (myininaya):

this will give you a positive result (as area should be positive or zero)

OpenStudy (anonymous):

Ah-ha! Thank you so much!

myininaya (myininaya):

np